OpenStudy (anonymous):

General factoring strategy to factor the polynomial x^2+10xy+25y^2-1

OpenStudy (vishweshshrimali5):

You can do this: \[\large{x^2 + 2*x*(5y) + (5y)^2 - 1}\] Can you write x^2 + 2*x*(5y) + (5y)^2 as some square ?

OpenStudy (vishweshshrimali5):

Remember : \[\large{a^2 + 2*a*b + b^2 = (a+b)^2}\]

hero (hero):

Use the difference of squares formula \(a^2 - b^2 = (a + b)(a - b)\) I can walk you through the steps if you like.